# Implementation of full and simplified likelihoods in CheckMATE Iñaki Lara and Krzysztof Rolbiecki

### v1

@@ 1 Introduction
In this paper, we discuss the implementation of the full and simplified likelihoods in the CheckMATE framework [14, 15, 16, 17]. For implementations in other tools, see eg. [18, 19]. In this paper, we include 9 ATLAS and 4 CMS searches. The CMS searches use the correlated background model (covariance matrix) [20]. For ATLAS searches, we use simplified [21] and full likelihood [22] frameworks. In addition, some of the implementations of the ATLAS searches include the full set of control regions. Users have several ways of controlling the actual calculation of likelihoods. The default method of evaluation is based on the package Spey [23]. We provide validation plots for each of the searches. A comparison of three methods for the calculation of limits is presented: the best-expected signal region, the full likelihood model (if available), and the simplified likelihood.
@@ 2.1 ATLAS
By default, the above calculation will be executed using the Spey program [23] and spey-pyhf plugin [29]. Spey is a Python-based cross-platform package that allows statistical inference of hypotheses using different likelihood prescriptions.♮$\natural$1♮$\natural$1$\natural$1Installation of Spey is straightforward: pip install spey. Please refer to Spey online documentation and spey-pyhf for more details [30]. In our setup it gives a somewhat better control over the calculation than the above-mentioned CheckMATE-pyhf interface, but nevertheless the calculation is still performed in the pyhf framework. The main motivation, however, for using Spey was a possibility of combining different searches (also between experiments), which is planned in the next release of CheckMATE. In any case, a direct evaluation using pyhf and bypassing Spey also remains available.
@@
[…]
@@ 2.2 CMS
cms_2205_09597 Search for production of charginos and neutralinos in final states containing hadronic decays of $WW$, $WZ$, or $WH$ and $E_{\mathrm{T}}^{\mathrm{miss}}$ 35 [42]
@@ 2.3 CheckMATE parameters
CheckMATE provides several switches and parameters to control details of statistical evaluation. These are summarized in Tab. 3. The switches are divided into two groups: one providing a control of what statistical tests are performed and the other to control different modes of calculation. By default no statistical evaluation is performed. For the sake of speed and stability one switch, scan, provides a quick and reliable way of obtaining Allowed/Excluded result but with limited additional information. Generally the available statistics include $\mathrm{CL}_{s}$ tests and calculation of upper limits on signal strength, both of which can be obtained as observed and/or expected measures. By choosing a select switch users can control which statistics are calculated. If no explicit choice is made the observed upper limit will be calculated. Finally, the detailed switch can be used to request calculation of all available statistics, but it should be noted that its execution can be time consuming. The option -so can be used to request calculation of statistics for previous CheckMATE runs (it requires presence of the evaluation/total_results.dat file in the output directory).
The second group of parameters is used to choose a method of calculation of requested statistics for the ATLAS searches (it does not affect the calculation for the CMS searches as described in the previous Section). For the default method, simple, calculation is performed using simplified likelihood and the CheckMATE-pyhf interface. The full switch chooses a calculation using the full likelihood and the CheckMATE-Spey interface. Finally, the fullpyhf switch requests calculation sing the full likelihood and CheckMATE-pyhf interface (this is somewhat less flexible regarding the output compared to the previous options). In any case, users should remember that the full likelihood calculation can be time consuming if many searches and signal regions are requested.
@@
Backend: X numpy Different backends for calculations in the full model.
@@
The results of the calculation for each of the multibin signal regions and all requested analyses are stored in the multibin_limits/results.dat file. In order to follow the progress of calculation the observed limits are also displayed on screen for each of the signal regions. The final evaluation is decided using the upper limit on the signal strength $\mu$ (if calculated):
@@ 2.4 Evaluation time
An example calculation was performed for the evaluation of the full likelihood model in the ATLAS search atlas_2101_01629, which is a moderately complex one of those implemented. The tests were performed on three Ubuntu 22.04.5 systems: A - Intel i7-4770 CPU 3.4 GHz, 4/8 cores/threads;♮$\natural$2♮$\natural$2$\natural$2It is really old. B - Intel i7-8565U CPU 1.8 GHz, 4/8 cores/threads and NVIDIA GeForce MX150; C - Intel i7-7700 CPU 3.4 GHz, 4/8 cores/threads. Jax, pytorch and tensorflow take advantage of the multithread computation even when no GPUs are present. On system B the CUDA 12.8 toolkit was installed for the GPU computation. In Table 4 we show the results. Note that the system B was tested in two configurations, with and without, GPUs. The best performance is obtained with jax and pytorch. For comparison, we also provide evaluation times for the $\mathrm{CL}_{s}$ observed limit performed on system A.
@@ 3 Validation
The validation procedure is organized as follows. For SUSY processes events are generated using MadGraph5_aMC@NLO 3.1.0 [47, 48, 49] with up to two additional partons in the final state. The NNPDF23LO [50, 51, 52] parton distribution function (PDF) set is used. The events are then interfaced to Pythia-8.2 [53, 54] for modeling of decays, hadronization and showering. The matrix element and parton shower matching was done using the CKKW-L [55] prescription and a matching scale of 1/4 of the SUSY particle mass was set. Inclusive signal cross sections for production of squarks and gluinos are obtained at the approximate next-to-next-to-leading order with soft gluon resummation at next-to-next-to-leading-logarithm (approximate NNLO+NNLL) [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66], following recommendations of Ref. [67]. The signal cross sections for production of sleptons, charginos and neutralinos are computed at next-to-leading order plus next-to-leading-log precision using Resummino [68, 69, 70, 71, 72, 73, 74]. This setup generally follows procedures employed within LHC experiments to obtain exclusion limits used in the validation process.
@@
We follow the same key for the plots throughout this section. The black solid (dashed) line is for ATLAS/CMS observed (expected) limit. On some occasions this is accompanied by theoretical (experimental) ±$\pm 1\sigma$ uncertainty. The red lines are used for the exclusion obtained with CheckMATE using the best signal region method, solid and dashed for observed and expected limits, respectively. Finally, the blue and green lines denote limits obtained using the simplified and full likelihoods, respectively.
@@
This is a search [32] for production of electroweakinos and sleptons. The final states with two leptons (opposite sign and same or different flavor SF/DF) and $E_{\mathrm{T}}^{\mathrm{miss}}$are considered. Events are first separated into SF and DF categories and further subdivided by the multiplicity of the non-b-tagged jets. A combined multi-bin SR is defined out of the 36 exclusive binned signal regions. The full likelihood statistical model is provided.
@@
The validation plots in Fig. 3 show a similar sensitivity of the best signal region, as well as full and simplified likelihoods. In each case, good agreement with the ATLAS result is observed. The observed exclusion limits slightly exceed the ATLAS limit, whereas for the expected limits, the best agreement is obtained for the full-likelihood approach, and comfortably inside $1\sigma$ experimental uncertainty. It should be noted that including the control region in the full likelihood fit was essential in obtaining the acceptable agreement.
@@
This is a search [34] for production of electroweakinos and sleptons in scenarios with compressed mass spectra. The final states contain two low $p_{T}$ leptons (opposite sign and same or different flavor). Sensitivity of the search relies on additional initial state radiation (ISR) jets which give transverse boost to the final state particles and adds missing transverse momentum. There are two multibin signal regions implemented in CheckMATE:♮$\natural$3♮$\natural$3$\natural$3The vector boson fusion (VBF) SRs are currently not available. SR-EWK targeting production of electroweakinos and divided into 44 bins according to the lepton pair invariant mass, $E_{\mathrm{T}}^{\mathrm{miss}}$, and lepton flavor; SR-S targeting production of sleptons and divided into 32 bins. The full likelihood model is provided.
@@
In Figure 5, we show validation plots for the wino-like chargino-neutralino pair production (EW search region). The searches for different parities of LSP and NLSP are shown separately: upper row $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$ (same parity) and $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$ (opposite parity) in the lower row. Electron and muon channels are shown separately. These examples show a danger of using a simplified likelihood model, which clearly produces too strong constraints.
@@
This is a search [36] for supersymmetric partners of top quark decaying to a Higgs or $Z$ boson. The final state Higgs boson is reconstructed from a pair of $b$-jets while the $Z$ boson from a same-flavor opposite-sign dilepton pair. There are 3 multibin signal regions that are shape fits in $E_{\mathrm{T}}^{\mathrm{miss}}$, $E_{\mathrm{T}}^{\mathrm{miss}}$-significance and $p_{\mathrm{T}}$ of the $Z$ candidate. No likelihood model is provided, and only simplified likelihood model is available.
@@
This is a search [37] for squarks (1st and 2nd generation) and gluinos in final states with 2-6 jets, 0 leptons and missing transverse momentum. There are three multibin signal regions in this search: MB-SSd which targets production of squarks and is divided into 24 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets; MB-GGd which targets production of gluinos and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$; MB-C which targets compressed spectra and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets. The control regions are also implemented what gives an opportunity to fully exploit the full likelihood model provided by the ATLAS collaboration.
Figure 8 shows validation plots for the squark pair-production process. The simplified shape-fit has excellent agreement with the official result, while the full likelihood model gives a somewhat weaker exclusion. In either case there is a clearly visible advantage over the best-SR method.
@@
As can be seen in Figure 10 we find excellent agreement between the result of the simplified shape-fit and the official result. The best-SR exclusion is significantly weaker. It should be noted that in the parts of the plot with large $m_{a}$ and/or $m_{A}$ the exclusion is driven by the shape-fit in the range of large $m_{T}$, which increases the sensitivity by factor $\sim 2$.
@@
For validation, we compare exclusion limits for the vector mediator model in the $m_{\textrm{mediator}}$$m_{\textrm{DM}}$ plane. The mediator-quarks and mediator-DM couplings are fixed to $g_{q}=0.25$ and $g_{\chi}=1$, respectively. The validation plot, Figure 13, shows again good agreement between the CheckMATE and CMS limits.
@@
This is a search [78] for charginos and neutralinos, that looks for final states with large missing transverse momentum and pairs of hadronically decaying bosons $WW$, $WZ$ and $WH$. This search makes use of specific algorithms (taggers) defined to identify $W$ boson/Higgs boson candidates out of the identified large-radius signal jets. The individual bins are defined orthogonally based on the number of $b$-tagged jets, the number of identified $W$, $Z$, and Higgs boson candidates and missing transverse momentum. The simplified likelihood is built out of 35 signal regions. Additionally, there are 4 aggregate signal regions defined.

### v2

@@ 1 Introduction
In this paper, we discuss the implementation of the full and simplified likelihoods in the CheckMATE framework [14, 15, 16, 17]. For implementations in other tools, see eg. [18, 19]. In this paper, we include 9 ATLAS and 4 CMS searches. The CMS searches use the correlated background model (covariance matrix) [20]. For ATLAS searches, we use the simplified [21] and the full likelihood [22] frameworks. In addition, some of the implementations of the ATLAS searches include the full set of control regions. Users have several ways of controlling the actual calculation of likelihoods. The default method of evaluation is based on the package Spey [23]. We provide validation plots for each of the searches. A comparison of three methods for the calculation of limits is presented: the best-expected signal region, the full likelihood model (if available), and the simplified likelihood.
@@ 2.1 ATLAS
[…]
@@
atlas_2111_08372 Search for associated production of a Z boson with an invisibly decaying Higgs boson or dark matter candidates $2 \usym1$F5F4 [39]
@@ 2.2 CMS
cms_2205_09597 Search for production of charginos and neutralinos in final states containing hadronic decays of $WW$, WZWZ, or $WH$ and $E_{\mathrm{T}}^{\mathrm{miss}}$ 35 [42]
@@ 2.3 CheckMATE parameters
CheckMATE provides several switches and parameters to control details of statistical evaluation. These are summarized in Tab. 3. The switches are divided into two groups: one providing a control of what statistical tests are performed and the other to control different modes of calculation. By default, no statistical evaluation is performed. For the sake of speed and stability, one switch, scan, provides a quick and reliable way of obtaining Allowed/Excluded result but with limited additional information. Generally, the available statistics include $\mathrm{CL}_{s}$ tests and calculation of upper limits on signal strength, both of which can be obtained as observed and/or expected measures. By choosing a select switch users can control which statistics are calculated. If no explicit choice is made, the observed upper limit will be calculated. Finally, the detailed switch can be used to request calculation of all available statistics, but it should be noted that its execution can be time consuming. The option -so can be used to request calculation of statistics for previous CheckMATE runs (it requires presence of the evaluation/total_results.dat file in the output directory).
The second group of parameters is used to choose a method of calculation of requested statistics for the ATLAS searches (it does not affect the calculation for the CMS searches as described in the previous Section). For the default method, simple, calculation is performed using simplified likelihood and the CheckMATE-Pyhf interface. The full switch chooses a calculation using the full likelihood and the CheckMATE-Spey interface. Finally, the fullpyhf switch requests the calculation using the full likelihood and CheckMATE-Pyhf interface (this is somewhat less flexible regarding the output compared to the previous options). In any case, users should remember that the full likelihood calculation can be time consuming if many searches and signal regions are requested.
@@
Backend: X numpy Different backends for calculations in the full model; Numpy is always available.
@@
The results of the calculation for each of the multibin signal regions and all requested analyses are stored in the multibin_limits/results.dat file. In order to follow the progress of the calculation, the observed limits are also displayed on screen for each of the signal regions. The final evaluation is decided using the upper limit on the signal strength $\mu$ (if calculated):
@@ 2.4 Evaluation time
[…]
@@ 3 Validation
The validation procedure is organized as follows. For SUSY processes, events are generated using MadGraph5_aMC@NLO 3.1.0 [47, 48, 49] with up to two additional partons in the final state. The NNPDF23LO [50, 51, 52] parton distribution function (PDF) set is used. The events are then interfaced to Pythia 8.3 [53, 54] for modeling of decays, hadronization and showering. The matrix element and parton shower matching for ATLAS searches was performed using the CKKW-L [55] prescription and a matching scale of 1/4 of the SUSY particle mass was set. For CMS searches, the MLM [56, 48] matching scheme was used. Inclusive signal cross sections for production of squarks and gluinos are obtained at the approximate next-to-next-to-leading order with soft gluon resummation at next-to-next-to-leading-logarithm (approximate NNLO+NNLL) [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67], following recommendations of Ref. [68]. The signal cross sections for production of sleptons, charginos and neutralinos are computed at next-to-leading order plus next-to-leading-log precision using Resummino [69, 70, 71, 72, 73, 74, 75]. This setup generally follows procedures employed within LHC experiments to obtain exclusion limits used in the validation process.
Events for the simplified scenario of DM production in the CMS search EXO-20-004 [42] were generated at NLO with MadGraph5_aMC@NLO 3.1.0 and the DMsimp model [76]. Up to 2 additional partons were included and the FxFx [77] jet matching scheme was used. Samples for the 2HDM+a model and validation of HIGG-2018-26 [39] were simulated internally in CheckMATE interfaced to MadGraph5_aMC@NLO 3.4.1 and hadronized with Pythia 8.2.
@@
We follow the same key for the plots throughout this section. The black solid (dashed) line is for ATLAS/CMS observed (expected) limit. On some occasions, this is accompanied by theoretical (experimental) $\pm 1\sigma$ uncertainty. The red lines are used for the exclusion obtained with CheckMATE using the best signal region method, solid and dashed for observed and expected limits, respectively. Finally, the blue and green lines denote the limits obtained using the simplified and full likelihoods, respectively.
@@
This is a search [32] for production of electroweakinos and sleptons. The final states with two leptons (opposite sign and same or different flavor SF/DF) and $E_{\mathrm{T}}^{\mathrm{miss}}$are considered. Events are first separated into SF and DF categories and are further subdivided by the multiplicity of the non-b-tagged jets. A combined multi-bin SR is defined out of the 36 exclusive binned signal regions. The full likelihood statistical model is provided.
@@
The validation plots in Fig. 3 show a similar sensitivity of the best signal region, as well as full and simplified likelihoods. In each case, good agreement with the ATLAS result is observed. The observed exclusion limits slightly exceed the ATLAS limit, whereas for the expected limits, the best agreement is obtained for the full-likelihood approach, and comfortably within the experimental $1\sigma$ uncertainty. It should be noted that including the control region in the full likelihood fit was essential in obtaining the acceptable agreement.
@@
[…]
@@
In Figure 5, we show validation plots for the wino-like chargino-neutralino pair production (EW search region). The searches for different parities of LSP and NLSP are shown separately: upper row $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$ (same parity) and $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$ (opposite parity) in the lower row; see eg. Ref. [78] for a discussion of CP properties of the neutralino sector and consequences for the decay kinematics. The electron and muon channels are shown separately. These examples show a danger of using a simplified likelihood model, which clearly produces too strong constraints.
@@
This is a search [36] for supersymmetric partners of top quark decaying to a Higgs or $Z$ boson. The final state Higgs boson is reconstructed from a pair of $b$-jets while the $Z$ boson from a same-flavor opposite-sign dilepton pair. There are 3 multibin signal regions that are shape fits in $E_{\mathrm{T}}^{\mathrm{miss}}$, $E_{\mathrm{T}}^{\mathrm{miss}}$-significance and $p_{\mathrm{T}}$ of the $Z$ candidate. No likelihood model is provided, and only a simplified likelihood model is available.
@@
This is a search [37] for squarks (1st and 2nd generation) and gluinos in final states with 2-6 jets, 0 leptons and missing transverse momentum. There are three multibin signal regions in this search: MB-SSd which targets production of squarks and is divided into 24 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets; MB-GGd which targets production of gluinos and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$; MB-C which targets compressed spectra and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets. The control regions are also implemented, which gives an opportunity to fully exploit the full likelihood model provided by the ATLAS Collaboration.
Figure 8 shows the validation plots for the squark pair-production process. The simplified shape-fit has excellent agreement with the official result, while the full likelihood model gives a somewhat weaker exclusion. In either case, there is a clearly visible advantage over the best-SR method.
@@
As can be seen in Figure 10 we find excellent agreement between the simplified shape fit result and the official result. The best-SR exclusion is significantly weaker. It should be noted that in the parts of the plot with large $m_{a}$ and/or $m_{A}$ the exclusion is driven by the shape fit in the range of large $m_{T}$, which increases the sensitivity by factor $\sim 2$.
@@
For validation, we compare exclusion limits for the vector mediator model in the $m_{\textrm{mediator}}$$m_{\textrm{DM}}$ plane. The mediator-quarks and mediator-DM couplings are fixed to $g_{q}=0.25$ and $g_{\chi}=1$, respectively. The validation plot, Figure 13, again shows good agreement between the CheckMATE and CMS limits.
@@
This is a search [82] for charginos and neutralinos, that looks for final states with large missing transverse momentum and pairs of hadronically decaying bosons $WW$, WZWZ and $WH$. This search makes use of specific algorithms (taggers) defined to identify $W$ boson/Higgs boson candidates out of the identified large-radius signal jets. The individual bins are defined orthogonally based on the number of $b$-tagged jets, the number of identified $W$, $Z$, and Higgs boson candidates and missing transverse momentum. The simplified likelihood is built out of 35 signal regions. Additionally, there are 4 aggregate signal regions defined.